OpenStudy (anonymous):

HELP I WILL MEDAL AND FAN PLZZZZZ ON A TIME LIMIT!!!!!!!!!!!!!!!!!! Solve. 18=5/6m A.21 3/5 B. 17 1/6 C. 15 D. 3/5

OpenStudy (anonymous):

do it urself figure it out xcant solve or will be suspende3de

OpenStudy (anonymous):

What is 18 divided by 5/6? That's the same as saying 18 multiplied by 6/5, right? 18/1 * 6/5 = 108/5. What do you get when you convert 108/5 to a mixed number?

OpenStudy (anonymous):

21 3/5

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

Yes, that's exactly right. You can check your work by multiplying 5/6 by 21 3/5. The result is 18.

OpenStudy (anonymous):

can some one help with a 2 more

OpenStudy (anonymous):

Which equation can be used to solve this problem Maura spent 1 1/2 hours on science homework. This was 2/3 of the total time she spent on homework How much time di she spend on homework A. h=2/3x1 1/2 B. 1 1/2 x h=2/3. C.h-2/3=1 1/2 D. 2/3h=1 1/2

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I will medal

OpenStudy (anonymous):

B

OpenStudy (retireed):

How many hours were spent on science homework?

OpenStudy (retireed):

S = 1.5 hours correct?

OpenStudy (retireed):

How much of the total homework time was spent on doing science?

OpenStudy (retireed):

science work was two thirds of the total homework time S = (2/3) H correct equation 2 ?

OpenStudy (anonymous):

so is it B

OpenStudy (retireed):

take equation 2 and substitute the S into it from the first equation S = 1.5

OpenStudy (anonymous):

I am lost

OpenStudy (retireed):

What is the second equation?

OpenStudy (anonymous):

sorry I am really bad at math one sec I have to do some thing be bake in like 5 min sorry brb

OpenStudy (anonymous):

ok one sec I will get the ?

OpenStudy (retireed):

S = 1.5 Eq.1 S = (2/3) H Eq.2 H is the total homework time

OpenStudy (anonymous):

Marley wanted each guest at her party to have a glass of lemonade. Each glass holds 8 1/3 ounces and there will be 7 guests how many ounces of lemonade will she for her guest A.58 1/3 oz B.56 1/3. oz C.25 2/3 oz D.15 1/3. oz

OpenStudy (retireed):

S = 1.5 S = (2/3) H Just wrote eq 1 slightly different 1.5 = S S = (2/3) H 1.5 = S <-- Plug in the value of S into ---> S = (2/3) H The answer is NOT B.